Consider a one-way ANOVA setup with three levels of an experimental factor, and with sample size n 15. Use R to obtain a randomization of treatments to experimental units, in such a way that someone else could duplicate your particular random assignment. Include your R code and output in your answer
Answer #1

### R command

Y=c(73,68,72,72,72,53,58,62,59,56,86,82,80,85,88)
Factor=c('1','1','1','1','1','2','2','2','2','2','3','3','3','3','3')
model=aov(Y~Factor)
summary(model)

## Output

> Y=c(73,68,72,72,72,53,58,62,59,56,86,82,80,85,88)
> Factor=c('1','1','1','1','1','2','2','2','2','2','3','3','3','3','3')
> model=aov(Y~Factor)
> summary(model)
Df Sum Sq Mean Sq F value Pr(>F)
Factor 2 1769.7 884.9 104.9 2.5e-08 ***
Residuals 12 101.2 8.4
---
Signif. codes: 0 ‘***’ 0.001 ‘**’ 0.01 ‘*’ 0.05 ‘.’ 0.1 ‘ ’ 1